Problem 7. A) Show that the potential energy of two particles is invariant under a Galilean transformation. Recall that U = U (t)-rı (t)) that is the separation is evaluated at a given time t. B) Show that the kinetic energy of an object is not, generally, invariant under a Galilean transformation. C) Argue that the total energy of a system of objects is not, generally, invariant under a Galilean transformation
Answer #1

(A)

U'=U'(vec r'_2(t)-vec r_1'(t))=U'(vec r_2(t)-vec v t-vec r_1(t)+vec v t)

=U(vec r_2(t)-vec r_1(t))

So it remains invariant.

(B)

r(t) = r(t)-vt

K'=rac {1}{2}mv'^2=rac {1}{2}m(v-V)^2 eq K=rac {1}{2}mv^2

(C)

Since the kinetic energy is not, generally, invariant under a Galilean transformation so the total energy is not, generally, invariant under a Galilean transformation.